OpenStudy (anonymous):

one qt of milk and two loaves of bread cost $3.54. three qts of milk and four loaves of bread cost $8.44. find the cost of each.

OpenStudy (radar):

I believe this has the making of a system of equations. Let b = the cost of a loaf of bread Let m= the cost of a quart of milk. Now we form the equations for our system. m + 2b = $3.54 that is the first statement in the problem. Now for the second. 3m + 4b = $8.44 Those two equations are what we need to solve. Do you know how to proceed to solve using either the substitution or the elimination method?

OpenStudy (anonymous):

No I get confused trying so solve them. I am trying to look up how to do them now but no luck.

OpenStudy (radar):

Usually they teach the "substitution" method first. So lets do it that way. we take one of the variables (b or m) and express it in terms of the other. In our problem, it is easier to substitute for m Using the first equation: m + 2b = 3.54 We express m in terms of b m = 3.54 - 2b Algebra, subtracted 2b from both sides of the equal sign. Now substitute for m in the second equation 3(3.54 - 2b) + 4b = 8.44 Now using algebra. 10.62 - 6b + 4b = 8.44 10.62 - 2b = 8.44 10.62 -8.44 = 2b 2.18 = 2b b= $1.09 Bread is $1.09 a loaf Now solve for m by substituting $1.09 for b in either one of the first equations and solve for the price for a quart of milk.

OpenStudy (radar):

Good luck with your studies.

OpenStudy (anonymous):

Thank you